2. An object is thrown upward and returns to the ground 12 seconds. How high did the ball go?

3. A bus is traveling with a uniform velocity of 20 m/s. At the instant it passes a car at rest, the car starts to accelerate at 4 m/s2 in the same direction as the velocity of the bus. (a) After how many seconds will the car overtake the bus? (b) How far will the car travel before overtaking the bus?
1
Expert's answer
2021-10-19T15:13:05-0400

2)

"h=\frac{g(\frac t2)^2}2=\frac{gt^2}8=176~m."

3)

"s=vt,"

"s=\frac{at^2}2,\implies"

(a)

"t=\frac{2v}a=10~s,"

(b)

"s=\frac{2v^2}a=200~m."